OpenStudy (anonymous):

Divide: 48xy4 - 72xy2 by 12xy2

OpenStudy (anonymous):

someone help

OpenStudy (kirbykirby):

\[ \frac{48xy^4-72xy^2}{12xy^2}=\frac{48xy^4}{12xy^2}-\frac{72xy^2}{12xy^2}\] Now, simplify the coefficients, and apply the exponent rules.. \[4 y^{4-2}- 6 \] The x's in the first fraction cancel out since it is of the same power in the numerator and denominator In the second fraction, all the x's and y'2 cancel out since they are of the same power in the numerator and denominator\

OpenStudy (anonymous):

thank you(:

OpenStudy (kirbykirby):

yw =]